Scot and Vidia, married taxpayers, earn $240,000 in taxable income and $5,000 in interest from an investment in City of Tampa bonds. (Use the U.S. tax rate schedule for married filing jointly).

2019 Tax Rate Schedules

Individuals

Schedule X-Single

If taxable income is over: But not over: The tax is:
$           0 $    9,700 10% of taxable income
$    9,700 $ 39,475 $970 plus 12% of the excess over $9,700
$ 39,475 $ 84,200 $4,543 plus 22% of the excess over $39,475
$ 84,200 $160,725 $14,382.50 plus 24% of the excess over $84,200
$160,725 $204,100 $32,748.50 plus 32% of the excess over $160,725
$204,100 $510,300 $46,628.50 plus 35% of the excess over $204,100
$510,300 $153,798.50 plus 37% of the excess over $510,300

Schedule Y-1-Married Filing Jointly or Qualifying Widow(er)

If taxable income is over: But not over: The tax is:
$           0 $ 19,400 10% of taxable income
$ 19,400 $ 78,950 $1,940 plus 12% of the excess over $19,400
$ 78,950 $168,400 $9,086 plus 22% of the excess over $78,950
$168,400 $321,450 $28,765 plus 24% of the excess over $168,400
$321,450 $408,200 $65,497 plus 32% of the excess over $321,450
$408,200 $612,350 $93,257 plus 35% of the excess over $408,200
$612,350 $164,709.50 plus 37% of the excess over $612,350

Schedule Z-Head of Household

If taxable income is over: But not over: The tax is:
$           0 $ 13,850 10% of taxable income
$ 13,850 $ 52,850 $1,385 plus 12% of the excess over $13,850
$ 52,850 $ 84,200 $6,065 plus 22% of the excess over $52,850
$ 84,200 $160,700 $12,962 plus 24% of the excess over $84,200
$160,700 $204,100 $31,322 plus 32% of the excess over $160,700
$204,100 $510,300 $45,210 plus 35% of the excess over $204,100
$510,300 $152,380 plus 37% of the excess over $510,300

Schedule Y-2-Married Filing Separately

If taxable income is over: But not over: The tax is:
$           0 $    9,700 10% of taxable income
$    9,700 $ 39,475 $970 plus 12% of the excess over $9,700
$ 39,475 $ 84,200 $4,543 plus 22% of the excess over $39,475
$ 84,200 $160,725 $14,382.50 plus 24% of the excess over $84,200
$160,725 $204,100 $32,748.50 plus 32% of the excess over $160,725
$204,100 $306,175 $46,628.50 plus 35% of the excess over $204,100
$306,175 $82,354.75 plus 37% of the excess over $306,175

Required:

  1. If Scot and Vidia earn an additional $80,000 of taxable income, what is their marginal tax rate on this income?
  2. What is their marginal rate if, instead, they report an additional $80,000 in deductions?

(For all requirements, round your answers to 2 decimal places.)

a.Marginal tax rate %

b.Marginal tax rate %
Answer #1

Answer-

Tax on $240000 = 28,765+(240000-168,400)*0.=24 45,949
a
Taxable income = 240000+80000 = $320000
Tax on $320000 = 28,765+(320000-168,400)*0.24= 65,149
Marginal tax rate = (65,149-45,949)/(320000-240000)= 24%
b
Taxable income = 240000-80000 = $160000
Tax on $160000 = 9,086+(160000-78,950)*0.22= 26,917
Marginal tax rate = (45,949-26,917)/(240000-160000)= 23.79%

Kindly give thumbs up if u like my answer...Thanks!!!